Can you please explain.
The probability is 1 in 4,000,000 that a single auto trip in the United States will result in a fatality. Over a lifetime, an average U.S. driver takes 50,000 trips. (a) What is the probability of a fatal accident over a lifetime? Explain your reasoning carefully. Hint: Assume independent events. Why might the assumption of independence be violated? (b) Why might a driver be tempted not to use a seat belt “just on this trip”?

Answer by checkley77(12844)

1/4,000,000=x/50,000
4,000,000x=50,000
x=50,000/4,000,000
x=.0125 chances of having an accident or 1.25%.
This assumes that the average length of the 4,000,000 & the 50,000 trips are about the same length/time.